# Mixtures of Neural Network Experts with Application to Phytoplankton Flow Cytometry Data

### v1

@@ Abstract
Flow cytometry is a valuable technique that measures the optical properties of particles at a single-cell resolution. When deployed in the ocean, flow cytometry allows oceanographers to study different types of photosynthetic microbes called phytoplankton. It is of great interest to study how phytoplankton properties change in response to environmental conditions. In our work, we develop a nonlinear mixture of experts model for simultaneous clustering and regression utilizing random-weight neural networks. Our model allows one to flexibly estimate how cell properties and relative abundances depend on environmental covariates, without the computational burden of backpropagation. We show that the proposed model provides superior predictive performance in simulated examples compared to a mixture of linear experts. Also, applying our model to real data, we show that our model has (1) comparable out-of-sample prediction performance, and (2) more realistic estimates of phytoplankton behavior.
@@ 1 Introduction
Phytoplankton are photosynthetic microbes that reside at the bottom of the marine food chain. They conduct more photosynthesis than all flora on land (Field et al.,, 1998), acting as an invaluable component of Earth’s biological carbon pump (Ducklow et al.,, 2015; Boyd et al.,, 2019). Thus, studying phytoplankton behavior and distribution in the ocean is directly relevant for our understanding of carbon transport and, by extension, the earth’s biogeochemical cycle (Mouw et al.,, 2016; Wang and Moore,, 2012). Phytoplankton have faced a number of environmental pressures in recent years, including increased levels of carbon dioxide, subsequent ocean acidification, and higher average annual temperatures (Van de Waal and Litchman,, 2020). Some recent work predicts major changes in relative abundances (Flombaum and Martiny,, 2021) in response to environmental changes. However, it is still largely unknown how these trends will affect phytoplankton abundance and photosynthetic activity. Thus, understanding the response of phytoplankton to various environmental conditions constitutes an important line of work.
@@
An increasingly popular method to measure phytoplankton attributes in the ocean is flow cytometry (Sosik et al.,, 2003). A fluid sample is passed through an instrument which measures the size and fluorescence for each particle. Then, the data analyst uses expert judgment – aided by software – to group these particles’ measurements into distinct types. This process is called gating and has traditionally been done manually, heavily relying on the analyst’s visual inspection of the data (Verschoor et al.,, 2015; BD Life Sciences,, 2023). In recent years, automatic gating algorithms have been in active development as an alternative to manual gating. However, most currently available gating algorithms are motivated by biomedical applications and are unsuitable for marine flow cytometry data. Our primary focus is to simultaneously cluster particles into distinct species, and to quantify each species’ response to changes in the ocean’s environmental conditions in a flexible manner, using a custom statistical mixture model.
@@
We briefly review the literature on statistical models for flow cytometry that are most relevant to our work. Recent model-based statistical methods typically assume the data come from a mixture of real-valued statistical distributions. Pyne et al., (2014) and Lee et al., (2016) use finite mixtures of multivariate skew $t$-distributions with sample-specific random effects. Dundar et al., (2014) and Cron et al., (2013) use hierarchical Dirichlet process mixtures of Gaussian kernels to borrow information across samples. Soriano and Ma, (2019) develop a specialized Dirichlet process Gaussian mixture which accounts for inter-sample variation in both the component parameters and weights. Gorsky et al., (2024) develop a Bayesian nonparametric mixture of hierarchical skew normal kernels, a hierarchy which captures both sample-specific and inter-sample structure. Hejblum et al., (2019) develop a Dirichlet process mixture of skew $t$-distributions which accounts for temporal ordering of samples. However, none of these approaches use environmental covariates. The most relevant method, outlined in Hyun et al., (2023), applies a mixture of experts (Jordan and Jacobs,, 1994) to phytoplankton flow cytometry data, modeling both cluster means and probabilities as linear functions of environmental covariates. We build on this last method by developing a mixture of nonlinear experts.
@@
The use of mixtures of experts to perform model-based clustering is supported by a rich literature which has developed over the past thirty years. Recently, Huang and Yao, (2012) showed that a semiparametric mixture of experts is identifiable under weak assumptions, and Nguyen et al., (2021) showed that a mixture of experts using a softmax gating function and location-scale kernels can provide arbitrarily accurate density estimates. Also, the use of a neural network in a mixture of experts has gained popularity in recent years. For example, Hadavandi et al., (2016) use a mixture of neural network experts in an agglomerative clustering scheme, and Liu et al., (2020) propose a mixture of neural network regressions in which all neural network experts share a common set of parameters. However, the latter approach does not use covariates to inform the cluster probabilities, and both are computationally intensive to train, requiring algorithms such as backprogagation or batch-size gradient descent. Our model also uses a mixture of neural network experts, but we randomly draw and then fix the hidden layer weights, thus bypassing the need for heavy computation. Huang et al., 2006a show that such random-weight neural networks can approximate any data-generating function with arbitrary accuracy and still generalize well to held-out samples. This technique has been applied to a variety of settings, including nonlinear regression (He et al.,, 2016), clustering (He et al.,, 2014; Ding et al.,, 2017), and mixture modeling (Husmeier and Taylor,, 1998). See Cao et al., (2018) for a more thorough review.
@@
The method developed by Hyun et al., (2023) assumes that Gaussian cluster means and the logits of the cluster probabilities are linear functions of the covariates. However, we show that the nonlinear functions are more apt for modeling flow cytometry data collected over a large spatial region. For instance, microbe characteristics such as cell size have biological upper and lower limits which linear functions cannot predict effectively. Generally, phytoplankton species’ attributes and abundance likely do not respond to environmental changes monotonically. Also, certain pairs of covariates may have an interaction effect – for example, increased nutrient availability are known to improve or hinder a cell’s ability to perform biological functions. Since the functional forms for nonlinear relationships are difficult to prescribe prior to modeling, a flexible nonparametric approach such as ours is useful.
@@
Here, we extend the approach of Hyun et al., (2023) in two ways. First, we use a random-weight neural network to expand the class of estimable regression functions. Such a model can more flexibly model cluster means and probabilities as functions of environmental conditions, thus allowing the user to fit a single model to flow cytometry data collected from a wide, heterogeneous range of ocean conditions. Second, we use the principal components of environmental covariates as regressors rather than the original covariates themselves. This vastly reduces the number of regressors and removes the strong intercorrelations in the original covariates, thus easing model interpretation.
@@ 2 SeaFlow cruise data description
We will treat marine flow cytometry data as a scatterplot-valued discrete time series $\{\boldsymbol{y}^{(t)}\}_{t=1}^{T}$, where each scatterplot $\boldsymbol{y}^{(t)}$ is an $n_{t}\times d$ matrix. We call $\boldsymbol{y}^{(t)}$ a cytogram. Each cytogram has rows $\boldsymbol{y}_{i}^{(t)}\in\mathbb{R}^{d}$ consisting of the $d$ measured optical properties of the $i$th particle observed during time interval $t$, for $i=1,\ldots,n_{t}$. The total number of particles observed is $N=\sum_{t=1}^{T}n_{t}$. Finally, cytogram $t$ shares covariate values $\boldsymbol{x}_{t}\in\mathbb{R}^{p}$, where $p\ll n_{t}$ for all $t$.
@@
The flow cytometry data we analyze are from the Gradients 2 cruise collected by the Armbrust Lab (Ribalet et al.,, 2019) in the summer of 2017. This cruise traversed a large spatial region with heterogeneous environmental conditions, making it a good candidate to showcase the flexibility of our model to variation in the relationships between response data and covariates. Specifically, the cruise traveled between the North Pacific Subpolar and Subtropical gyres, collecting measurements of phytoplankton abundance and optical properties using the continuous-time flow cytometer known as SeaFlow (Swalwell et al.,, 2011). Each particle $\boldsymbol{y}_{i}^{(t)}$ from a given hour-long interval $t$ has $d=3$ measurements: (1) cell diameter, and (2) red fluorescence and (3) orange fluorescence, corresponding to the light intensities of emission spectra associated with chlorophyll and phycoerythrin pigments. Since all three dimensions’ original measurements are highly skewed, they are transformed using the natural logarithm (then linearly rescaled to be roughly between 0 and 8). In total, there are $083$-dimensional cytograms. However, we later created a twelve-hour lagged version of a covariate, which led us to remove the first 12 cytograms for a final count of 296. Additionally, some observations were censored at SeaFlow’s measurement limits; these points were simply removed prior to modeling. Lastly, an auxiliary measurement $\{c_{i}^{(t)}\}_{i,t}$, the inferred biomass of each particle, will be used. We use this cytogram and biomass data as a regression response.
@@
Accompanying the response data, there are $33$ environmental covariates sourced from the Simons CMAP database (Ashkezari et al.,, 2021), accessed via the cmap4r R package (Hyun et al.,, 2025), and aggregated to match the hourly time resolution of the response data. These environmental covariate data include biochemical and physical attributes collected from satellites (and processed as long-term averages) and from on-board instruments, for a total of $p=37$ covariates over $T=296$ time points corresponding to unique hours. The vector of all $37$ covariate values at time $t$ is written as $\boldsymbol{x}_{t}$, and $\boldsymbol{X}\in\mathbb{R}^{296\times 37}$ is the covariate matrix with row $t$ equal to $\boldsymbol{x}_{t}$. For more details regarding data curation, see Hyun et al., (2023).
@@
Figure 1: Principal components of the environmental covariates. The top panel shows the sample correlations between the first four principal components, latitude, and the covariates. The bottom-left panel shows the first principal component (PC1), as well as the most scientifically relevant covariates which are strongly correlated with PC1, plotted over time. We also overlay latitude as a dashed black line to show its similarity to PC1. The two vertical dotted lines at latitudes 31.9 N and 34 N represent the estimated boundaries of the North Pacific Transition Zone, a region where latitude-varying conditions change rapidly, hand-picked as the region where PC1 changes rapidly. In the bottom-right panel, we display a similar plot for PC3 and PC4. The full set of principal components is shown in Supplement A.
The 37 covariates are numerous and strongly intercorrelated, so we apply principal components analysis (PCA) to the covariate matrix. We first calculate the singular value decomposition of the covariate matrix $\boldsymbol{X}$ as
$\boldsymbol{X}=\boldsymbol{U}\boldsymbol{D}\boldsymbol{V}^{\prime}$
and then calculate the $q$-dimensional orthogonal projection of the covariate matrix $\boldsymbol{\Psi}\in\mathbb{R}^{T\times q}$ as
$\boldsymbol{\Psi}=\boldsymbol{X}\boldsymbol{V}_{\cdot,1:q},$ (1)
@@
where $\boldsymbol{V}_{\cdot,1:q}$ are the first $q$ columns of $\boldsymbol{V}$. In all experiments, we choose $q$ to be the lowest value such that the proportion of variance explained by the first $q$ principal components is at least 0.95. For the dataset used in our analysis, we found $q=9$, which we use throughout this article.
@@
It is worth examining the principal components $\{\boldsymbol{\Psi}_{\cdot,j}\}_{j=1}^{4}$ in some detail. For brevity we will adopt the shorthand of, for example, PC1 to denote the first principal component. As shown in Figure 1, the first principal component PC1 has a strong correlation with latitude. Since PC1 and latitude share strong associations with environmental conditions such as nutrient levels, temperature, and salinity, information about these conditions can be compressed into a single variable which summarizes latitudinal variation in the ocean. The second principal component PC2 is most strongly associated with wind speed, sea level anomaly, and iron levels, summarizing information relating to physical oceanographic characteristics. PC3 is related to physical ocean characteristics such as zonal geostrophic velocities and lagged sunlight (photosynthetically active radiation or PAR for short) time-lagged by 3 and 6 hours. Lastly, PC4 is comprised of PAR and its 9- and 12- hour lagged variants.
@@ 3 Methodology
We assume cytograms follow a mixture of multivariate Gaussian distributions
@@
with observations independent across all particles $i$ and all time intervals $t$. Next, we model the mixture component means and probabilities as a regression on environmental covariates. However, instead of directly regressing on the principal components of the covariates, we regress on the hidden layer outputs of a random-weight neural network (Huang et al., 2006b, ):
@@
Here, $\boldsymbol{\beta}_{0,k}\in\mathbb{R}^{d\times 1}$, $\boldsymbol{\beta}_{k}\in\mathbb{R}^{n_{h}\times d}$, $\alpha_{0,k}\in\mathbb{R}^{1}$, and $\boldsymbol{\alpha}_{k}\in\mathbb{R}^{n_{h}}$ are regression coefficients. Also, $\sigma\left(\boldsymbol{h}_{t}\right)$ is a random-weight neural network with $n_{h}$ hidden nodes, constructed as follows. We first create a matrix $\boldsymbol{W}\in\mathbb{R}^{(p+1)\times n_{h}}$ of hidden layer weights and biases by randomly and independently drawing each element from a $\text{Unif}\left(-a,a\right)$ distribution, for a fixed constant $a$. Then, the principal components are transformed as
$\boldsymbol{h}_{t}=\boldsymbol{W}^{\prime}\begin{pmatrix}1\\ \boldsymbol{\Psi}_{t,\cdot}\end{pmatrix},$ (2)
where $\boldsymbol{\Psi}_{t,\cdot}$ is a column vector taken from row $t$ of $\boldsymbol{\Psi}$. Next, we apply a nonlinear activation function to extract nonlinear features. Specifically, denoting the elements of $\boldsymbol{h}_{t}$ as $h_{t,1},h_{t,2},\ldots,h_{t,n_{h}}$, we apply the softmax function (the element-wise logistic function):
@@
Crucially, the hidden weights are randomly drawn once and are fixed before the model training process, so the computational cost of estimating hidden layer weights via backpropagation is entirely avoided. Since a random-weight neural network can approximate any data-generating function with arbitrary accuracy (Huang et al., 2006a, ), it allows our model to learn nonlinear relationships between the response data’s mixture model parameters and the environmental covariates.
The approach taken in Hyun et al., (2023) is a special case of this model when $\boldsymbol{V}_{1:q,\cdot}=\boldsymbol{I}_{p}$, $\boldsymbol{W}=\boldsymbol{I}_{p+1}$, and the activation function $\sigma(\cdot)$ is set to be the identity function $\sigma(\boldsymbol{h}_{t})=\boldsymbol{h}_{t}$. In Sections 4 and 5 we compare our proposed model to one we call “linear”, which is identical to the one developed by Hyun et al., (2023) but uses principal components as regressors. This allows us to easily compare the interpretations of different models – nonlinear and linear – estimated using a common set of regressors. Also, using $q=9$ principal components enables simpler interpretation than if we had used the original $p=37$ covariates.

### v2

@@ Abstract
Flow cytometry is a valuable technique that measures the optical properties of particles at a single-cell resolution. When deployed in the ocean, flow cytometry allows oceanographers to study different types of photosynthetic microbes called phytoplankton. It is of great interest to study how phytoplankton properties change in response to environmental conditions. In our work, we develop a nonlinear mixture of experts model to estimate separate regression functions for each subpopulation utilizing random-weight neural networks. Our model allows one to flexibly estimate how cell properties and relative abundances depend on environmental covariates in each segment of a heterogeneous sample, without the computational burden of backpropagation. We show that the proposed model provides superior predictive performance in simulated examples compared to a mixture of linear experts. Also, applying our model to real data, we show that our model has (1) comparable out-of-sample prediction performance, and (2) more realistic estimates of phytoplankton behavior.
@@ 1 Introduction
Phytoplankton are photosynthetic microbes that reside at the bottom of the marine food chain. They conduct more photosynthesis than all flora on land [11], acting as an invaluable component of Earth’s biological carbon pump [8, 3]. Thus, studying phytoplankton behavior and distribution in the ocean is directly relevant for our understanding of carbon transport and, by extension, the earth’s biogeochemical cycle [33, 45]. Phytoplankton have faced a number of environmental pressures in recent years, including increased levels of carbon dioxide, subsequent ocean acidification, and higher average annual temperatures [43]. Some recent work predicts major changes in relative abundances [12] in response to environmental changes. However, it is still largely unknown how these trends will affect phytoplankton abundance and photosynthetic activity. Thus, understanding the response of phytoplankton to various environmental conditions constitutes an important line of work.
@@
An increasingly popular method to measure phytoplankton attributes in the ocean is flow cytometry [39]. A fluid sample is passed through an instrument which measures the size and fluorescence for each particle. Then, the data analyst uses expert judgment – aided by software – to group these particles’ measurements into distinct types. This process is called gating and has traditionally been done manually, heavily relying on the analyst’s visual inspection of the data [44, 2]. In recent years, automatic gating algorithms have been in active development as an alternative to manual gating. However, most currently available gating algorithms are motivated by biomedical applications and are unsuitable for marine flow cytometry data. Our primary focus is to quantify each species’ response to changes in the ocean’s environmental conditions in a flexible manner, using a custom statistical mixture model. We adopt the mixture of experts framework to separate particles into subpopulations, fit a separate regression on each subpopulation, and also fit a regression on each cluster probability. This allows a simple framework for domain experts to understand the relationships between all subpopulation properties and environmental conditions.
@@
We briefly review the literature on statistical models for flow cytometry that are most relevant to our work. Recent model-based statistical methods typically assume the data come from a mixture of real-valued statistical distributions. [35] and [28] use finite mixtures of multivariate skew $t$-distributions with sample-specific random effects. [9] and [5] use hierarchical Dirichlet process mixtures of Gaussian kernels to borrow information across samples. [38] develop a specialized Dirichlet process Gaussian mixture which accounts for inter-sample variation in both the component parameters and weights. [14] develop a Bayesian nonparametric mixture of hierarchical skew normal kernels, a hierarchy which captures both sample-specific and inter-sample structure. [19] develop a Dirichlet process mixture of skew $t$-distributions which accounts for temporal ordering of samples. Specific to oceanography, [25] use an expectation-maximization (EM) algorithm to estimate a trend-filtered mixture of Gaussians which automatically gates cells into subpopulations. [6] build on this technique by providing a more computationally lightweight solution via kernel smoothing. However, none of these approaches use environmental covariates. Thus, these models are suitable for clustering a population into subpopulations, not for inference on the response of each subpopulation to environmental changes, which is our primary interest. The most relevant method, outlined in [24], applies a mixture of experts [27] to phytoplankton flow cytometry data, modeling both cluster means and probabilities as linear functions of environmental covariates. We build on this last method by developing a mixture of nonlinear experts.
@@
The use of mixtures of experts to jointly perform regression and clustering is supported by a rich literature which has developed over the past thirty years. Recently, [22] showed that a semiparametric mixture of experts is identifiable under weak assumptions, and [34] showed that a mixture of experts using a softmax gating function and location-scale kernels can provide arbitrarily accurate density estimates. Also, the use of a neural network in a mixture of experts has gained popularity in recent years. For example, [16] use a mixture of neural network experts in an agglomerative clustering scheme, and [29] propose a mixture of neural network regressions in which all neural network experts share a common set of parameters. However, the latter approach does not use covariates to inform the cluster probabilities, and both are computationally intensive to train, requiring algorithms such as backprogagation or batch-size gradient descent. Our model also uses a mixture of neural network experts, but we randomly draw and then fix the hidden layer weights, thus bypassing the need for heavy computation. [20] show that such random-weight neural networks can approximate any continuous data-generating function with arbitrary accuracy and still generalize well to held-out samples. This technique has been applied to a variety of settings, including nonlinear regression [18], clustering [17, 7], and mixture modeling [23]. See [4] for a more thorough review.
@@
The method developed by [24] assumes that Gaussian cluster means and the logits of the cluster probabilities are linear functions of the covariates. However, we show that the nonlinear functions are more apt for modeling flow cytometry data collected over a large spatial region. For instance, microbe characteristics such as cell size have biological upper and lower limits which linear functions cannot predict effectively. Generally, phytoplankton species’ attributes and abundance likely do not respond to environmental changes monotonically. Also, certain pairs of covariates may have an interaction effect – for example, increased nutrient availability are known to improve or hinder a cell’s ability to perform biological functions. Apart from these general hypotheses, very little is known about the functional forms of these nonlinear relationships. This lack of prior knowledge makes our flexible nonparametric approach to the regression functions useful.
@@
Here, we extend the approach of [24] in two ways. First, we use a random-weight neural network to expand the class of estimable regression functions. Such a model can more flexibly model cluster means and probabilities as functions of environmental conditions, thus allowing the user to fit a single model to flow cytometry data collected from a wide, heterogeneous range of ocean conditions. Second, we use the principal components of environmental covariates as regressors rather than the original covariates themselves. This vastly reduces the number of regressors and removes the strong intercorrelations in the original covariates, thus easing model interpretation.
@@ 2 SeaFlow cruise data description
The flow cytometry data we analyze are from the Gradients 2 cruise collected by the Armbrust Lab [36] in the summer of 2017. This cruise traversed a large spatial region with heterogeneous environmental conditions, making it a good candidate to showcase the flexibility of our model to variation in the relationships between response data and covariates. Specifically, the cruise traveled between the North Pacific Subpolar and Subtropical gyres, collecting measurements of phytoplankton abundance and optical properties using the continuous-time flow cytometer known as SeaFlow [40]. We refer to each observation collected by SeaFlow as a particle, since it is a collection of optical properties of a particle suspended in ocean water. We assume particles are aggregated to discrete time, so each time point $t$ contains multiple ($n_{t}$) particles collected over a wider time interval (in our case, an hour), for $t=1,\ldots,T$. We denote the $i$th particle at time $t$ as $\boldsymbol{y}_{i}^{(t)}\in\mathbb{R}^{d}$, for $i=1,\ldots,n_{t}$ and refer to the collection of all particles at time $t$ as a cytogram, denoted by $\boldsymbol{y}^{(t)}$. SeaFlow measures $d=3$ properties of each particle: (1) cell diameter, and (2) red fluorescence and (3) orange fluorescence, corresponding to the light intensities of emission spectra associated with chlorophyll and phycoerythrin pigments. Since all three dimensions’ original measurements are highly skewed, they are transformed using the natural logarithm, resulting in roughly Gaussian cluster shapes. We then linearly rescale the data to be roughly between 0 and 8 in each dimension. Imposing a common scale is useful for visualization and for specifying direction-invariant constraints on cluster movement (see (4)).
In total, there are $308$ 3-dimensional cytograms, each consisting of all particles collected during one hour of the cruise. However, we later created a twelve-hour lagged version of a covariate, which led us to remove the first 12 cytograms for a final count of $T=296$. Additionally, some observations were censored at SeaFlow’s measurement limits; these points were simply removed prior to modeling. Lastly, an auxiliary measurement $\{c_{i}^{(t)}\}_{i,t}$, the inferred biomass of each particle, will be used. We use this cytogram and biomass data as a regression response.
@@
Accompanying the response data, there are $33$ environmental covariates sourced from the Simons CMAP database [1], accessed via the cmap4r R package [32], and aggregated to match the hourly time resolution of the response data. These environmental covariate data include biochemical and physical attributes collected from satellites (and processed as long-term averages) and from on-board instruments, for a total of $p=37$ covariates over $T=296$ time points corresponding to unique hours. The vector of all $37$ covariate values at time $t$ is written as $\boldsymbol{x}_{t}$, and $\boldsymbol{X}\in\mathbb{R}^{296\times 37}$ is the covariate matrix with row $t$ equal to $\boldsymbol{x}_{t}$. For more details regarding data curation, see [24].
@@
Figure 1: Principal components of the environmental covariates. The top panel shows the sample correlations between the first four principal components, latitude, and the covariates. The bottom-left panel shows the first principal component (PC1), as well as the most scientifically relevant covariates which are strongly correlated with PC1, plotted over time. We also overlay latitude as a dashed black line to show its similarity to PC1. The two vertical dotted lines at latitudes 31.9 N and 34 N represent the estimated boundaries of the North Pacific Transition Zone, a region where latitude-varying conditions change rapidly, hand-picked as the region where PC1 changes rapidly. In the bottom-right panel, we display a similar plot for PC3 and PC4. In Supplement A, we display a similar plot for PC2 (Supplementary Figure S1) and the full set of principal components (Supplementary Figure S2).
@@
The 37 covariates are numerous and strongly intercorrelated, so we apply principal components analysis (PCA) to the covariate matrix. The resulting design matrix has columns consisting of the first $q$ principal components and is denoted by $\boldsymbol{\Psi}\in\mathbb{R}^{T\times q}$. In all experiments, we choose $q$ to be the lowest value such that the proportion of variance explained by the first $q$ principal components is at least 0.95. For the dataset used in our analysis, we found $q=9$, which we use throughout this article. We show in Supplement F that model interpretation does not change much for values of $q$ larger than 9.
@@
It is worth examining the principal components $\{\boldsymbol{\Psi}_{\cdot,j}\}_{j=1}^{4}$ in some detail. For brevity we will adopt the shorthand of, for example, PC1 to denote the first principal component. As shown in Figure 1, the first principal component PC1 has a strong correlation with latitude. Since PC1 and latitude share strong associations with environmental conditions such as nutrient levels, temperature, and salinity, information about these conditions can be compressed into a single variable which summarizes latitudinal variation in the ocean. The second principal component PC2 is most strongly associated with wind speed, sea level anomaly, and iron levels, summarizing information relating to physical oceanographic characteristics. PC3 is related to physical ocean characteristics such as zonal geostrophic velocities and lagged sunlight (photosynthetically active radiation or PAR for short) time-lagged by 3 and 6 hours. Lastly, PC4 is comprised of PAR and its 9- and 12- hour lagged variants. Visual inspection of PCs 5–9 revealed no easily discernible biological interpretations, so we omit them from subsequent analysis.
@@ 3 Methodology
Although the cytograms $\left\{\boldsymbol{y}^{(t)}\right\}_{t=1}^{T}$ are collected sequentially, it is reasonable from an oceanographic standpoint to assume that spatio-temporal variation in microbial communities is largely due to heterogeneity in environmental conditions. In comparison, variation due to time and space alone is negligible. Therefore, we assume the cytograms are conditionally exchangeable given the covariates. That is, for any permutation $\tau$ of $\{1,\cdots,T\}$, the probabilistic model for the particles is of the form
$p\left(\boldsymbol{y}^{(\tau(1))},\cdots,\boldsymbol{y}^{(\tau(T))}\mid\boldsymbol{x}_{\tau(1)},\ldots,\boldsymbol{x}_{\tau(T)}\right)=p\left(\boldsymbol{y}^{(1)},\cdots,\boldsymbol{y}^{(T)}\mid\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{T}\right),$
where $p$ is the joint density of the cytograms induced by a probabilistic model for the particles. In our mixture of experts, we assume each cytogram follows a mixture of multivariate Gaussian distributions
@@
with observations independent across all particles $i$ and all times $t$. Next, we model the mixture component means and probabilities as a regression on environmental covariates. However, instead of directly regressing on the principal components of the covariates, we regress on the hidden layer outputs of a random-weight neural network [21]:
@@
Here, $\boldsymbol{\beta}_{0,k}\in\mathbb{R}^{d\times 1}$, $\boldsymbol{\beta}_{k}\in\mathbb{R}^{n_{h}\times d}$, $\alpha_{0,k}\in\mathbb{R}^{1}$, and $\boldsymbol{\alpha}_{k}\in\mathbb{R}^{n_{h}}$ are regression coefficients. Also, $\sigma\left(\boldsymbol{h}_{t}\right)$ is a random-weight neural network with $n_{h}$ hidden nodes, constructed as follows. We draw the $m$th hidden node’s weights and biases $w_{m,1},\cdots,w_{m,p},b_{m}\overset{\text{iid}}{\sim}\text{Unif}(-a,a)$, for a fixed constant $a$. This is repeated for $m=1,\cdots,n_{h}$, where $n_{h}$ is the total number of hidden nodes. The distribution used to generate the weights and biases (collectively called hidden weights) should not be interpreted as a prior distribution, since the hidden weights are not updated at all during the model training process. After they are drawn, they are considered fixed, known constants. Then, the principal components are transformed as
$h_{t,m}=b_{m}+\sum_{j=1}^{p}w_{m,j}\psi_{t,j}.$ (1)
That is, $h_{t,m}$ is a random linear combination of the principal components at time $t$. Next, we apply a nonlinear activation function to extract nonlinear features. Specifically, collecting the hidden nodes into a vector $\boldsymbol{h}_{t}=(h_{t,1},h_{t,2},\ldots,h_{t,n_{h}})^{\prime}$, we apply the logistic function element-wise:
@@
The vector $\boldsymbol{h}_{t}$ can also be represented more succinctly as
$\boldsymbol{h}_{t}=\boldsymbol{W}\begin{pmatrix}1\\ \boldsymbol{\Psi}_{t,\cdot}\end{pmatrix}$,
where $\boldsymbol{W}\in\mathbb{R}^{n_{h}\times(p+1)}$ is a matrix containing the hidden weights.
Crucially, the hidden weights are randomly drawn once and are fixed before the model training process, so the computational cost of estimating hidden layer weights via backpropagation is entirely avoided. Since a random-weight neural network can approximate any data-generating function with arbitrary accuracy [20], it allows our model to learn nonlinear relationships between the response data’s mixture model parameters and the environmental covariates. The element-wise logistic function is not a strict requirement; the same guarantees hold for any element-wise infinitely differentiable function. We chose the logistic function due to its ubiquity in both random and non-random neural network literature.
The approach taken in [24] is a special case of this model when $\boldsymbol{V}_{1:q,\cdot}=\boldsymbol{I}_{p}$, $n_{h}=p$, $\boldsymbol{W}=\boldsymbol{I}_{p+1}$, and the activation function $\sigma(\cdot)$ is set to be the identity function $\sigma(\boldsymbol{h}_{t})=\boldsymbol{h}_{t}$. In Sections 4 and 5 we compare our proposed model to one we call “linear”, which is identical to the one developed by [24] but uses principal components as regressors. This allows us to easily compare the interpretations of different models – nonlinear and linear – estimated using a common set of regressors. Also, using $q=9$ principal components enables simpler interpretation than if we had used the original $p=37$ covariates.